Kurz Co is an all equity firm with 20 million shares outstanding and has a stock price of $8 per share. Although investors currently expect Kurz to remain an all equity firm, Kurz plans to announce that it will borrow 50 million and use the funds to repurchases shares. Kurz will pay interest only and never plans to increase or decrease the debt amount. Kurz pays a 40% tax rate.
A)   What is the market value of Kurz’s existing assets before the announcement?
B)   What is the market value of Kurz’s assets (including any tax shields) just after the debt is issued, but before the shares are repurchased?
C)   What is Kurz’s share price just before the share repurchases and share price after the repurchases?
D)   What is Kurz’s total firm value after the shares are repurchased?
Answer #1

Answer A) Assets = Equities + Debt .. but debt is not given before the announcement. Thus, Assets = Equities
Market Value of Assets before the announcement = 20 million shares * $8 per share = $160 million

Answer B) after the debt is issued
Market Value of Assets = $160 mn + $50mn (cash) + Tax shield
= $210 mn + 40% of 50 = $230 million

Answer C) Kurz balance sheet before purchases
Debt + Equity = Assets
Equity = Assets - Debt = $230mn -$50mn = $ 180mn
Thus, Share Price before purchases = 180mn /20 mn shares = $9 per share

Kurz balance sheet after purchases
Assets = $160mn (existing) + tax shield = 160 + 20 = $180mn
Debt = $50mn
thus, Equity =$130mn
Therefore, Number of shares after repurchase = 20,000,000 - (50,000,000/9) = 20,000,000 - 5,555,555.55 shares  
  = 14,444,444.44 shares
Price of share after purchase = $130 mn / 14,444,444.44 share =$9 per share

Answer D) Kurz's firm value after shares are purchased
Assets = $180mn